OpenStudy (anonymous):

an isosceles triangle has a vertical angle of 108 degree and a base 20m long. Calculate its altitude. plz reply me fast

OpenStudy (anonymous):

Do 108 + 20 x 2 I think

OpenStudy (anonymous):

no it is from trigonometry

OpenStudy (mathmath333):



OpenStudy (anonymous):

Well at least he got 100

OpenStudy (anonymous):

no it is not the answer, the answer is 7.27cm

OpenStudy (anonymous):

this question is from D3 Ex:10f Q:2

OpenStudy (mathmath333):

u imagine a base of 20 meters and height of just 7.27 cm????

OpenStudy (mathmath333):

ok wait

OpenStudy (mathmath333):

oh my god u r right

OpenStudy (anonymous):

Lol

OpenStudy (mathmath333):

\(\large\tt \color{black}{tan54=\dfrac{10}{x}}\) \(\large\tt \color{black}{x=\dfrac{10}{tan54}}\) \(\large\tt \color{black}{x=\dfrac{10}{1.37}}\) \(\large\tt \color{black}{x=7.26~~m}\)

OpenStudy (mathmath333):

and its unit will be in meter

OpenStudy (mathmath333):

sry lol

OpenStudy (anonymous):

for what?

OpenStudy (mathmath333):

\(\Huge\tt \color{black}{\ddot\smile}\)

OpenStudy (anonymous):

I meant he got 100 percent on his test

OpenStudy (anonymous):

no this is not test it is assiement